# MonoBox: Tightness-free Box-supervised Polyp Segmentation using Monotonicity Constraint

### v4

@@ Abstract
Keywords:
Polyp segmentation Weak supervision Noise Label.
footnotetext: Equal Contribution; Corresponding Authors.
1 Introduction
@@
Colorectal Cancer (CRC) threats to human health worldwide, and colonoscopy is a golden-standard of identifying and resecting early polyps [12]. Recently, numerous deep learning-based polyp segmentation methods [8, 42, 43, 7] have been proposed to assist the accurate resection and workload reduction. However, they are mostly fully-supervised and thus require pixel-level mask labels, which are time-consuming and expensive to acquire.
Figure 1: (a) Examples of real non-tight box annotations in the in-house dataset. The red dashed lines indicate regions where the annotation is too wide, and the yellow dashed lines indicate regions where the annotation is too narrow. (b) In cases of four typical noisy box annotations (green line), our proposed monotonicity constraint (MC) provides correct constrains for four sampled bags (black dashed line) from the unconfident region (the region between two yellow lines), but the traditional MIL lead to incorrect constraints.
@@
To reduce the annotation cost, weakly-supervised segmentation (WSS) methods are studied, aiming to train segmentation models with more cost-effective labels, such as image-level categories [2], points [4], and bounding boxes [31, 33]. Among them, box-supervised segmentation (BSS) methods achieve the closest performance to fully-supervised methods, and thus attract dominant research attentions. The prevailing idea of BSS is called multiple-instance learning (MIL), which views each pixel as an instance, and defines a pixel-width image column or row as a positive bag if it crosses the annotated box, or negative bag otherwise. By pooling the instances’ predictions as the corresponding bag-level prediction, the segmentation model can be trained to produce pixel-level results in the optimization of bag classification. However, the existing MIL-based BSS methods [31, 33, 5, 37, 36] mostly, if not all, are based on a box-tightness assumption, that is, the box’s edges must precisely touch the target to make sure the bag-level labels are accurate. This severely inhibits the practical application of polyp datasets. In the polyp context, as shown in Fig. 1(a), the characteristics of polyps, such as blurred boundaries, low contrast with normal tissues, and small sizes, bring ambiguities to annotators and lead to inaccurate (i.e., non-tight) box annotations. Therefore, improving the tolerance of BSS methods to inaccurate boxes can alleviate the annotation difficulty and thus is of urgency for the clinical usage of polyp segmentation.
@@
The above objective can fall into the scope of noise learning, where the noise is induced by box non-tightness. A straightforward solution is to rectify the boxes to enhance the tightness [29, 41], but this usually requires a set of clean data as reference, which is beyond the scope of this work. Without clean reference, one promising solution is to apply noise-tolerant classification losses [35, 21] for addressing the inaccurate bag-level labels in MIL. In addition to the loss term, sampling more accurate bags is another direction [34, 44]. However, no matter improving classification losses or sampling strategies in MIL, these solutions ignore the spatial correlation of noises (i.e., incorrect bags), showing weakness in the task of segmentation.
@@
In this paper, we propose MonoBox that constrains monotonicity for tightness-free BSS, which is motivated by two characteristics of the spatial distribution of incorrect bags: (1) the confusing bags are sampled nearby the non-tight box edges, and (2) the probability of sampling a positive bag decrease from inside to outside across the box edges. Specifically, we first define the regions near the box edges as unconfident regions and the others as confident ones. For bags sampled from the confident regions, we adopt traditional MIL constraints. For bags sampled from the unconfident regions, we propose a novel monotonicity constraint (MC), which does not force hard predictions aligning the unreliable labels, but instead encourages a monotonicity trend that the inner response should be higher than the outer response. Principally, MC finds a soft but reliable surrogate objective when the precise ground-truth is absent, and the objective conforms to the expected spatial distribution pattern. As evidenced by the four typical noisy cases in unconfident regions (see Fig. 1(b)), for the bags across the inaccurate box edge (i.e., a,b,c,d), the MIL’s classification loss may result in misguiding supervisions, while MC can always derive meaningful gradients for optimization.
@@
2 Related Work
2.1 Fully-supervised Polyp Segmentation
@@
The polyp segmentation methods mostly rely on the fully-supervised learning to train their models. For example, SFA [9] explicitly applied an area-boundary constraint to supervise the learning of both polyp regions and boundaries. PraNet [8] generated a global map as the initial guidance region and then used the reverse attention module to refine the segmentation results by using the multilevel features. SANet [38] adopted a shallow attention module and a color exchange operation to remove background noise and reduce the interference of image color to segmentation respectively. Polyp-PVT [7] utilized a transformer encoder, a cascaded fusion module, a camouflage identification module, and a similarity aggregation module, to effectively suppress noises in the features and significantly improve their expressive capabilities. However, these methods all require pixel-level polyp annotations, which is difficult and time-consuming to acquire than box-level annotations, impeding the application on large-scale datasets.
2.2 Weakly-supervised Segmentation using Boxes
@@
For reducing the cost of annotation, box-supervised segmentation is extensively studied. These methods can be broadly classified into two categories. One is generating pseudo-labels and the other is based on multiple-instance learning (MIL).
@@
Earlier, most of the methods were based on generating pseudo-labels. For example, BoxSup [6], Box2Seg [18], and Boxshrink [11] used non-learning techniques, MCG [39], GrabCut [27], and SLIC [1], respectively generates refined pixel-level masks as pseudo labels based on ground truth (GT) boxes. Mahani et al. [23] proposed to reduce the loss weight of high entropy regions according to the predicted results to reduce error propagation during training. Oh et al. [24] refined pseudo-labels according to homogeneity of the backgournd inside and outside the GT box.
@@
Recently, an idea based on MIL has been adopted by a wide range of methods due to its effectiveness [15, 31, 19, 5, 36]. It takes the compactness of the box as a priori and considers the maximum value of any row or column of pixels within the box as a positive and the maximum value of any row or column of pixels not passing through the box as a negative. The implementation details vary from method to method, such as projection into a 1D vector [31, 19, 5] or reconstruction into a 2D box-like mask [37, 36]. Besides the MIL loss, some auxiliary constraints, such as color parwise affinity [31], multi-scale consistency [37] are also included to further improve the performance. Among them, IBoxCLA [36] proposed Contrastive Latent-Anchors (CLA), which enhances the feature contrast between the polyp and the surrounding normal tissue, and achieves state-of-the-art (SOTA) in box-supervised polyp segmentation. However, these methods are extremely dependent on the box tightness assumption, which reduces the cost-effective advantage of box-supervised methods.
2.3 Learning with Noisy Labels
@@
The task that training accurate models using noisy labels has been an active research area. In the classification realm, various techniques were proposed, such as label correction [22, 29], noise-tolerant loss function [10, 21], and data cleaning [13, 17]. Most of them can be translated to segmentation task which can be regarded as pixel-level classification. In the detection realm, He et al. [14] proposed KL loss for learning localization variance to alleviate the interference of ambiguity boxes on the detector. Xu et al. [41] introduced a meta-learning method to deal with noisy labels by utilizing a few clean samples. OA-MIL and SSD-Det [40] proposed to leverage clean class labels as guidance signals for refining inaccurate bounding boxes. However, there are drawbacks when applying to segmentation. The solutions for classification do not consider the structure information of the object, while those for detection can not well guarantee the tightness in corrected boxes.
@@
At present, there are only a few studies explicitly aiming at box-supervised segmentation with noisy box labels. PolarT [34] adopt MIL on the polar transformed image, which reduces the number of incorrect bags but biases the model toward simple instances. FSRM [44] first generated pseudo-masks based on the noisy boxes, and then relied on the pseudo-masks to guide the bag sampling in MIL. However, the noises also can be inherited in the pseudo-masks, negatively impacting the following MIL learning consequently. In this paper, we aim to seek a noise-tolerant constraint to replace the unreliable MIL-loss on the non-tight regions, and view the constraint as a plug-and-play module to boost the robustness of the current MIL-based SOTAs to tightness-free box annotations.
3 Method
@@
Fig. 2 gives the overview of MonoBox, which employs the most advanced and efficient image-level MIL fashion [37, 36] by optimizing a proxy map rather than individual sampled bags for box-supervised segmentation, and introduces a new monotonicity constraint (MC) to improve the optimization reliability on the noisy regions. Meanwhile, MonoBox designs a label correction strategy to dynamically improve the tightness of box annotations. In the following, we first recap the optimization of proxy map, and then explain the key components of MonoBox, and at last briefly discuss the extensibility of MC to other box-supervised MIL frameworks of segmentation.
3.1 Optimizing Proxy Map for Box-supervised Segmentation
@@
Given a colonoscopy image $I\in\mathbbm{R}^{3\times H\times W}$ and its GT box annotation $B=(x_{lt},y_{lt},x_{rb},y_{rb})$, where $(x_{lt},y_{lt})$ and $(x_{rb},y_{rb})$ represent the coordinates of the top-left and bottom-right points of box, respectively, the GT box-filled mask $b\in{\{0,1\}}^{H\times W}$ is created by assigning 1 within $B$ and 0 outside $B$, and a segmented map $m\in(0,1)^{H\times W}$ is obtained by applying the segmentation model on $I$.
@@
where $i$ and $j$ represent the pixel indexes. The proxy map $p$ decouples the shape information from the segmentation map, and thus can be optimized directly using the GT box-filled mask, as illustrated in Fig. 2. The typical optimization objective is the consistency constraint (e.g., Dice) between $b$ and $p$, which is formulated as follows:
@@
Figure 2: Overview of our proposed MonoBox. For the proxy map decoupled from the segmented map, we first define the confident and unconfident regions, and then adopt the consistency constraint and monotonicity constraint on them, respectively. Moreover, we utilize a strategy of label correction strategy to dynamically improve the tightness of box annotations
3.2 Separating Confident and Unconfident Regions
Based on the fact that the noise in the $m$ caused by non-tight box distributes around the edges, we divide the proxy map into confident and unconfident regions. Specifically, there are four unconfident regions $\{R_{u}^{l},R_{u}^{r},R_{u}^{t},R_{u}^{b}\}\in\mathbbm{R}^{1\times H\times W}$, corresponding to the left, right, top, and bottom edges of the GT box, respectively. The areas of four unconfident regions are determined by the width and height of the GT box annotation and an unconfident scale $\lambda$. For example, $R_{u}^{l}$ can be calculated as:
$\small R_{u}^{l}\left({x,y}\right)=\left\{\begin{array}[]{lr}{1,~{}~{}x\in\left[{x_{lt}-\lambda\cdot w,x_{lt}+\lambda\cdot w}\right]~{}and~{}y\in\left[{y_{lt}-\lambda\cdot h,y_{rb}+\lambda\cdot h}\right],}&\\ {0,~{}~{}otherwise.}\end{array}\right.$ (3)
@@ 3.2 Separating Confident and Unconfident Regions
$R_{c}=\mathbbm{1}-(R_{u}^{l}\cup R_{u}^{r}\cup R_{u}^{t}\cup R_{u}^{b}),$ (4)
where $\mathbbm{1}$ is a map of size $H\times W$ with all ones. On the confident region, we adopt the consistency constraint, which is calculated in Eq. 2. On the unconfident region, we use our proposed monotonicity constraint, which is detailed in the next part.
3.3 Monotonicity Constraint on Unconfident Regions
@@
$f_{x^{-}}^{{}^{\prime}}=\begin{bmatrix}1&{-1}\\ 1&{-1}\end{bmatrix},f_{x^{+}}^{{}^{\prime}}=\begin{bmatrix}{-1}&1\\ {-1}&1\end{bmatrix},f_{y^{-}}^{{}^{\prime}}=\begin{bmatrix}1&1\\ {-1}&{-1}\end{bmatrix},f_{y^{+}}^{{}^{\prime}}=\begin{bmatrix}{-1}&{-1}\\ 1&1\end{bmatrix},$ (5)
and then we take $f_{x^{-}}^{{}^{\prime}}$, $f_{x^{+}}^{{}^{\prime}}$, $f_{y^{-}}^{{}^{\prime}}$, and $f_{y^{+}}^{{}^{\prime}}$ to compute the first-order gradient maps on the proxy map $p$ for $R_{u}^{l}$, $R_{u}^{r}$, $R_{u}^{t}$, and $R_{u}^{b}$, respectively. $p_{x^{-}}^{{}^{\prime}}$, $p_{x^{+}}^{{}^{\prime}}$, $p_{y^{-}}^{{}^{\prime}}$, and $p_{y^{+}}^{{}^{\prime}}$ are the four corresponding gradient maps, and utilized to calculate losses by comparing them with zeros for monotonicity constraint (MC loss). Since the constraint of zero may be too stringent, we utilize a small margin $\epsilon$ that allows limited outliers. Taking the MC loss in $R_{u}^{l}$ as an example, the formula is as follows:
$\mathcal{L}_{MC}^{l}={\sum\limits_{i,j|R_{u}^{l}{[i,j]}=1}{max\left(p_{x^{-}}^{{}^{\prime}}[i,j]-\epsilon,0\right).}}$ (6)
@@
3.4 Label Correction for Dynamic Training
@@
3.5 Generality for MIL-based Box-supervised Segmentation
Although the implementation of the monotonicity constraint is based on proxy map optimization, we would like to remark that it is easy to expand to other MIL-based variants [15, 31] of box-supervised segmentation. Specifically, for the bags sampled from the unconfident regions, we only need to remember the row or column index of each bag, and the monotonicity constraint is a contrastive constraint of bag pairs with adjacent indexes, that is, the prediction of bag with inside index should be higher than that with outside index. Such process is just simplified by use of the proxy map and the four derivative kernels in our implementation of MonoBox.
4 Experiments
4.1 Implementation Details
MonoBox is implemented using PyTorch [25] and trained using a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. We use AdamW [20] as the optimizer, and set both the learning rate and weight decay to 0.0001. We resize the input image into 352 × 352 and set the batch size to 16. We train the models for 50 epochs in total and evoke the label correction every 10 epochs. The unconfident scale $\lambda$ in Eq. 3, the small margin $\epsilon$ in Eq. 6, and the IoU threshold $\tau$ in Sec. 3.4 are set to 0.2, 0.1, and 0.7, respectively.
4.2 Datasets and Evaluation Metrics
4.2.1 Public Synthetic Noisy Dataset.
We select five public polyp datasets used in previous work [36]: ClinicDB [3], Kvasir-SEG [16], ColonDB [30], EndoScene [32], and ETIS [28]. Following [36], we set 550 samples in ClinicDB and 900 samples in Kvasir as the train-set, and the remaining samples from these two datasets and all samples from the other three datasets as the test-set. Note that the original annotations for these datasets are ground truth (GT) masks, and we convert the GT masks to boxes by finding the tightest bounding boxes of the connected components
@@ 4.2 Datasets and Evaluation Metrics
4.2.2 In-house Real Noisy Dataset.
The in-house dataset cosists 18,656 colonoscopy images of polyp. The dataset is from a local hospital, which is split to the train-set (17,350 images) and test-set (1,306 images). The train-set is provided with box annotations (Fig. 1 (a) shows four box-annotated examples), and the test-set is provided with pixel-level annotations by two experts. Written informed consent was not required for this study as documented clinical colonoscopic images were collected retrospectively and appropriately by anonymizing and deidentifying. Therefore, the study design was exempted from full review by the Institutional Review Board.
4.2.3 Evaluation Metrics.
@@
4.3 Comparison with Anti-noise SOTAs

### v5

@@ Abstract
Introduction
@@
Colorectal Cancer (CRC) threats to human health worldwide, and colonoscopy is a golden-standard of identifying and resecting early polyps (Haggar and Boushey 2009; Ji et al. 2024). Recently, numerous deep learning-based polyp segmentation methods (Fan et al. 2020; Zhao, Zhang, and Lu 2021; Dong et al. 2021; Zhang et al. 2022; Ji et al. 2023) have been proposed to assist the accurate resection and workload reduction. However, they are mostly fully-supervised and thus require pixel-level mask labels, which are time-consuming and expensive to acquire.
Figure 1: Examples of non-tight box annotations produced by endoscopists in the real annotation process. The red dashed lines indicate regions where the annotation is too wide, and the yellow dashed lines indicate regions where the annotation is too narrow.
@@
To reduce the annotation cost, weakly-supervised segmentation (WSS) methods are studied, aiming to train segmentation models with more cost-effective labels, such as image-level categories (Ahn, Cho, and Kwak 2019), points (Cheng, Parkhi, and Kirillov 2022), and bounding boxes (Tian et al. 2021; Wang and Xia 2021). Among them, box-supervised segmentation (BSS) methods achieve the closest performance to fully-supervised methods, and thus attract dominant research attentions. The prevailing idea of BSS is called multiple-instance learning (MIL), which views each pixel as an instance, and defines a pixel-width image column or row as a positive bag if it crosses the annotated box, or negative bag otherwise. By pooling the instances’ predictions as the corresponding bag-level prediction, the segmentation model can be trained to produce pixel-level results in the optimization of bag classification. However, the existing MIL-based BSS methods (Tian et al. 2021; Wang and Xia 2021; Cheng et al. 2023; Wei et al. 2023; Wang et al. 2023) mostly, if not all, are based on a box-tightness assumption, that is, the box’s edges must precisely touch the target to make sure the bag-level labels are accurate. This severely inhibits the practical application of polyp datasets. In the polyp context, as shown in Figure 1, the characteristics of polyps, such as blurred boundaries, low contrast with normal tissues, and small sizes, bring ambiguities to annotators and lead to inaccurate (i.e., non-tight) box annotations. Therefore, improving the tolerance of BSS methods to inaccurate boxes can alleviate the annotation difficulty and thus is of urgency for the clinical usage of polyp segmentation.
[…]
@@
The above objective can fall into the scope of noise learning, where the noise is induced by box non-tightness. A straightforward solution is to rectify the boxes to enhance the tightness (Song, Kim, and Lee 2019; Xu et al. 2021), but this usually requires a set of clean data as reference, which is beyond the scope of this work. Without clean reference, one promising solution is to apply noise-tolerant classification losses (Wang et al. 2019; Ma et al. 2020) for addressing the inaccurate bag-level labels in MIL. In addition to the loss term, sampling more accurate bags is another direction (Wang and Xia 2022; Zhu et al. 2023). However, no matter improving classification losses or sampling strategies in MIL, these solutions ignore the spatial correlation of noises (i.e., incorrect bags), showing weakness in the task of segmentation.
@@
In this paper, we propose MonoBox that constrains monotonicity for tightness-free BSS, which is motivated by two characteristics of the spatial distribution of incorrect bags: (1) the confusing bags are sampled nearby the non-tight box edges, and (2) the probability of sampling a positive bag decrease from inside to outside across the box edges. Specifically, we first define the regions near the box edges as unconfident regions and the others as confident ones. For bags sampled from the confident regions, we adopt traditional MIL constraints. For bags sampled from the unconfident regions, we propose a novel monotonicity constraint (MC), which does not force hard predictions aligning the unreliable labels, but instead encourages a monotonicity trend that the inner response should be higher than the outer response. Principally, MC finds a soft but reliable surrogate objective when the precise ground-truth is absent, and the objective conforms to the expected spatial distribution pattern. As evidenced by the four typical noisy cases in unconfident regions (see Figure 2, for the bags (i.e., a,b,c,d) across the inaccurate box edge, the MIL’s classification loss may result in misguiding supervisions, while MC can always derive meaningful gradients for optimization.
@@
Related Work
Fully-supervised Polyp Segmentation
@@
The polyp segmentation methods mostly rely on the fully-supervised learning to train their models. For example, PraNet (Fan et al. 2020) generated a global map as the initial guidance region and then used the reverse attention module to refine the segmentation results by using the multilevel features. SANet (Wei et al. 2021) adopted a shallow attention module and a color exchange operation to remove background noise and reduce the interference of image color to segmentation respectively. Polyp-PVT (Dong et al. 2021) utilized a transformer encoder, a camouflage identification module, and a similarity aggregation module, to effectively suppress noises in the features and significantly improve their expressive capabilities. However, these methods all require pixel-level polyp annotations, which is difficult and time-consuming to acquire than box-level annotations, impeding the application on large-scale datasets.
Weakly-supervised Segmentation using Boxes
@@
For reducing the cost of annotation, box-supervised segmentation (BSS) is extensively studied. These methods can be broadly classified into two categories. One is generating pseudo-labels and the other is based on multiple-instance learning (MIL).
@@
Earlier, most of the methods were based on generating pseudo-labels. For example, BoxSup (Dai, He, and Sun 2015) and Box2Seg (Kulharia et al. 2020) used MCG (Williams et al. 2004) and GrabCut (Rother, Kolmogorov, and Blake 2004), respectively to generate refined pixel-level masks as pseudo labels based on ground truth (GT) boxes. Mahani et al. (Mahani et al. 2022) proposed to reduce the loss weight of high entropy regions according to the predicted results to reduce error propagation during training. However, these methods rely heavily on the quality of pseudo labels and lack a stable and accurate guidance.
@@
Recently, an idea based on MIL has been adopted by a wide range of methods due to its effectiveness (Hsu et al. 2019; Tian et al. 2021; Lan et al. 2023; Cheng et al. 2023; Wang et al. 2023). It takes the compactness of the box as a priori and considers the maximum value of any row or column of pixels within the box as a positive and the maximum value of any row or column of pixels not passing through the box as a negative. The implementation details vary from method to method, such as projection into a 1D vector (Tian et al. 2021; Lan et al. 2023; Cheng et al. 2023) or reconstruction into a 2D box-like mask (Wei et al. 2023; Wang et al. 2023). With MIL loss as the pipeline, some auxiliary constraints, such as color parwise affinity (Tian et al. 2021), multi-scale consistency (Wei et al. 2023) are introduced to further improve the performance. Among them, IBoxCLA (Wang et al. 2023) proposed Contrastive Latent-Anchors (CLA), which enhances the feature contrast between the polyp and the surrounding normal tissue, and achieves state-of-the-art (SOTA) in box-supervised polyp segmentation. However, these methods are extremely dependent on the box tightness assumption, which reduces the cost-effective advantage of BSS methods.
Learning with Noisy Labels
@@
The task that training accurate models using noisy labels has been an active research area. In the classification realm, various techniques were proposed, such as label correction (Ma et al. 2018; Song, Kim, and Lee 2019), noise-tolerant loss function (Ghosh, Kumar, and Sastry 2017; Ma et al. 2020), and data cleaning (Han et al. 2018; Jiang et al. 2018). Most of them can be translated to segmentation task which can be regarded as pixel-level classification. In the detection realm, He et al. (He et al. 2019) proposed KL loss for learning localization variance to alleviate the interference of ambiguity boxes on the detector. Xu et al. (Xu et al. 2021) introduced a meta-learning method to deal with noisy labels by utilizing a few clean samples. OA-MIL and SSD-Det (Wu et al. 2023) proposed to leverage clean class labels as guidance signals for refining inaccurate bounding boxes. However, there are drawbacks when applying to segmentation. The solutions for classification do not consider the structure information of the object, while those for detection can not well guarantee the tightness in corrected boxes.
@@
At present, there are only a few studies explicitly aiming at BSS with noisy box labels. PolarT (Wang and Xia 2022) adopt MIL on the polar transformed image, which reduces the number of incorrect bags but biases the model toward simple instances. FSRM (Zhu et al. 2023) first generated pseudo-masks based on the noisy boxes, and then relied on the pseudo-masks to guide the bag sampling in MIL. However, the noises also can be inherited in the pseudo-masks, negatively impacting the following MIL learning consequently. In this paper, we aim to seek a noise-tolerant constraint to replace the unreliable MIL-loss on the non-tight regions, and view the constraint as a plug-and-play module to boost the robustness of the current MIL-based SOTAs to tightness-free box annotations.
Method
@@
Figure 3(b) gives the overview of MonoBox, which employs the most advanced and efficient image-level MIL fashion (Wei et al. 2023; Wang et al. 2023) by optimizing a proxy map rather than individual sampled bags for box-supervised segmentation (BSS), and introduces a new monotonicity constraint (MC) to improve the optimization reliability on the noisy regions. Meanwhile, MonoBox designs a label correction strategy to dynamically improve the tightness of box annotations. In the following, we first recap the optimization of proxy map, and then explain the key components of MonoBox, and at last briefly discuss the extensibility of MC to other box-supervised MIL frameworks of segmentation.
Optimizing Proxy Map for BSS
@@
Given a colonoscopy image $I\in\mathbb{R}^{3\times H\times W}$ and its GT box annotation $B=(x_{lt},y_{lt},x_{rb},y_{rb})$, where $(x_{lt},y_{lt})$ and $(x_{rb},y_{rb})$ represent the coordinates of the top-left and bottom-right points of box, respectively, the GT box-filled mask $b\in{\{0,1\}}^{H\times W}$ is created by assigning 1 within $B$ and 0 outside $B$, and a segmented map $m\in(0,1)^{H\times W}$ is obtained by applying the segmentation model on $I$.
@@
where $i$ and $j$ represent the pixel indexes. The proxy map $p$ decouples the shape information from the segmentation map, and thus can be optimized directly using the GT box-filled mask, as illustrated in Figure 3(b). The typical optimization objective is the consistency constraint (e.g., Dice) between $b$ and $p$, which is formulated as follows:
@@
Figure 3: (a) Schema of generating GT box-filled mask (assigning 1 within box annotations and 0 outside box annotations), confident region (as Eq. 4) and unconfident regions (as Eq. 3). (b) Overview of our proposed MonoBox. For the proxy map decoupled from the segmented map, we first define the confident and unconfident regions, and then adopt the consistency constraint and monotonicity constraint on them, respectively. Moreover, we utilize a strategy of label correction strategy to dynamically improve the tightness of box annotations
Separating Confident and Unconfident Regions
Based on the fact that the noise in the $m$ caused by non-tight box distributes around the edges, we divide the proxy map into confident and unconfident regions. Specifically, as shown in Figure 3(a), there are four unconfident regions $\{R_{u}^{l},R_{u}^{r},R_{u}^{t},R_{u}^{b}\}\in\mathbb{R}^{1\times H\times W}$, corresponding to the left, right, top, and bottom edges of the GT box, respectively. The areas of four unconfident regions are determined by the width and height of the GT box annotation and an unconfident scale $\lambda$. For example, $R_{u}^{l}$ can be calculated as:
$R_{u}^{l}\left({x,y}\right)=\left\{\begin{array}[]{lr}{1,~{}~{}x\in\left[{x_{lt}-\lambda\cdot w,x_{lt}+\lambda\cdot w}\right]}&\\ {~{}~{}~{}~{}~{}and~{}y\in\left[{y_{lt}-\lambda\cdot h,y_{rb}+\lambda\cdot h}\right],}&\\ {0,~{}~{}otherwise.}\end{array}\right.$ (3)
@@
$R_{c}=\mathbf{J}-(R_{u}^{l}\cup R_{u}^{r}\cup R_{u}^{t}\cup R_{u}^{b}),$ (4)
where $\mathbf{J}$ is a map of size $H\times W$ with all ones. On the confident region, we adopt the consistency constraint, which is calculated in Eq. 2. On the unconfident region, we use our proposed monotonicity constraint, which is detailed in the next part.
Monotonicity Constraint on Unconfident Regions
@@
$f_{x^{-}}^{{}^{\prime}}=\begin{bmatrix}1\\ -1\end{bmatrix}^{T}f_{x^{+}}^{{}^{\prime}}=\begin{bmatrix}{-1}\\ {1}\end{bmatrix}^{T}f_{y^{-}}^{{}^{\prime}}=\begin{bmatrix}1\\ {-1}\end{bmatrix}f_{y^{+}}^{{}^{\prime}}=\begin{bmatrix}{-1}\\ 1\end{bmatrix},$ (5)
and then we take $f_{x^{-}}^{{}^{\prime}}$, $f_{x^{+}}^{{}^{\prime}}$, $f_{y^{-}}^{{}^{\prime}}$, and $f_{y^{+}}^{{}^{\prime}}$ to compute the first-order gradient maps on the proxy map $p$ for $R_{u}^{l}$, $R_{u}^{r}$, $R_{u}^{t}$, and $R_{u}^{b}$, respectively. $p_{x^{-}}^{{}^{\prime}}$, $p_{x^{+}}^{{}^{\prime}}$, $p_{y^{-}}^{{}^{\prime}}$, and $p_{y^{+}}^{{}^{\prime}}$ are the four corresponding gradient maps, and utilized to calculate losses by comparing them with zeros for monotonicity constraint (MC loss). Taking the MC loss in $R_{u}^{l}$ as an example, the formula is as follows:
$\mathcal{L}_{MC}^{l}={\sum\limits_{i,j|R_{u}^{l}{[i,j]}=1}{max\left(p_{x^{-}}^{{}^{\prime}}[i,j],0\right).}}$ (6)
@@
Label Correction for Dynamic Training
@@
Generality for MIL-based BSS
Although the implementation of the MC is based on proxy map optimization, we would like to remark that it is easy to expand to other MIL-based variants (Hsu et al. 2019; Tian et al. 2021) of BSS. Specifically, for the bags sampled from the unconfident regions, we only need to remember the row or column index of each bag, and the monotonicity constraint is a contrastive constraint of bag pairs with adjacent indexes, that is, the prediction of bag with inside index should be higher than that with outside index. Such process is just simplified by use of the proxy map and the four derivative kernels in our implementation of MonoBox.
Experiments
Implementation Details
MonoBox is implemented using PyTorch (Paszke et al. 2019) and trained using a single NVIDIA GeForce RTX 3090 GPU with 24GB memory. We use AdamW (Loshchilov and Hutter 2017) as the optimizer, and set both the learning rate and weight decay to 0.0001. We resize the input image into 352 × 352 and set the batch size to 16. We train the models for 50 epochs in total and evoke the label correction every 10 epochs, i.e., $t=10$. The unconfident scale $\lambda$ in Eq. 3 and the IoU threshold $\tau$ in label correction are set to 0.2 and 0.7, respectively.
Datasets and Evaluation Metrics
Public Synthetic Noisy Dataset.
We select five public polyp datasets used in previous work (Wang et al. 2023): ClinicDB (Bernal et al. 2015), Kvasir-SEG (Jha et al. 2020), ColonDB (Tajbakhsh, Gurudu, and Liang 2015), EndoScene (Vázquez et al. 2017), and ETIS (Silva et al. 2014). Following (Wang et al. 2023), we set 550 samples in ClinicDB and 900 samples in Kvasir as the train-set, and the remaining samples from these two datasets and all samples from the other three datasets as the test-set. Note that the original annotations for these datasets are ground truth (GT) masks, and we convert the GT masks to boxes by finding the tightest bounding boxes of the connected components.
@@
In-house Real Noisy Dataset.
The in-house dataset cosists 18,656 colonoscopy images of polyp. The dataset is from a local hospital, which is split to the train-set (17,350 images) and test-set (1,306 images). The train-set is provided with box annotations (Figure 1 shows four box-annotated examples), and the test-set is provided with pixel-level annotations by two experts. Written informed consent was not required for this study as documented clinical colonoscopic images were collected retrospectively and appropriately by anonymizing and deidentifying. Therefore, the study design was exempted from full review by the Institutional Review Board.
Evaluation Metrics.
@@
Comparison with Anti-noise SOTAs